# Predictability Enables Parallelization of Nonlinear State Space Models

### v2

@@ Abstract
The rise of parallel computing hardware has made it increasingly important to understand which nonlinear state space models can be efficiently parallelized. Recent advances like DEER [1] or DeepPCR [2] have shown that evaluating a state space model can be recast as solving a parallelizable optimization problem, and sometimes this approach can yield dramatic speed-ups in evaluation time. However, the factors that govern the difficulty of these optimization problems remain unclear, limiting the larger adoption of the technique. In this work, we establish a precise relationship between the dynamics of a nonlinear system and the conditioning of its corresponding optimization formulation. We show that the predictability of a system, defined as the degree to which small perturbations in state influence future behavior, impacts the number of optimization steps required for evaluation. In predictable systems, the state trajectory can be computed in $\mathcal{O}((\log T)^{2})$ time, where $T$ is the sequence length, a major improvement over the conventional sequential approach. In contrast, chaotic or unpredictable systems exhibit poor conditioning, with the consequence that parallel evaluation converges too slowly to be useful. Importantly, our theoretical analysis demonstrates that for predictable systems, the optimization problem is always well-conditioned, whereas for unpredictable systems, the conditioning degrades exponentially as a function of the sequence length. We validate our claims through extensive experiments, providing practical guidance on when nonlinear dynamical systems can be efficiently parallelized, and highlighting predictability as a key design principle for parallelizable models.
@@ 1 Introduction
via the Gauss-Newton method. [^1] DEER [1] and DeepPCR [2] were concurrent works that both proposed to use the Gauss-Newton method for optimizing nonlinear sum of squares to parallelize sequential processes. In this paper, we therefore use DEER, DeepPCR, and Gauss-Newton interchangeably. Gonzalez et al. [3] further developed these methods, including quasi-Newton methods and trust regions methods for parallel evaluation of nonlinear dynamical systems. These methods evaluate nonlinear dynamical systems by iteratively linearizing the nonlinear system and evaluating the resulting linear dynamical system (LDS) with a parallel (a.k.a. associative) scan [4, 5]. Each parallel evaluation of an LDS implements one optimization step  [1, 2, 3, 6].
@@ 2 Problem Statement & Background
We define the residual and corresponding merit [^2] While minimizing a “merit function” is admittedly counterintuitive, we follow Nocedal and Wright [19, see eq. 11.35] in this convention. function $\mathcal{L}$ by stacking the elements $s_{t}\in\mathbb{R}^{D}$ of a trajectory into a $TD$-dimensional vector $\mathbf{s}$ and considering the vector of temporal differences,
@@ 3.1 The Merit Function is PL
3.2 Merit Function PL Constant is Controlled by the Largest Lyapunov Exponent of Model
@@ 4 Rates of Convergence for Optimizing the Merit Function
See Appendix H. We make no claim about the originality of lower bounding the size of the basin of quadratic convergence in Gauss-Newton. In fact, our proof of ˜5 closely follows the convergence analysis of Newton’s method in Section 9.5.3 of Boyd and Vandenberghe [28]. Our contribution is we highlight the elegant way the predictability $\lambda$ and non-linearity $L$ of a dynamical system influence an important feature of its merit function’s landscape. ∎
@@ 5 Experiments
We conduct experiments to support the theory developed above, demonstrating that predictability enables parallelization of nonlinear SSMs. To illustrate this point, we use Gauss-Newton optimization (aka DEER [1, 2]). We provide more experimental details in Appendix J. Our code is at https://github.com/lindermanlab/predictability_enables_parallelization
@@
Theorem 2 predicts a sharp phase transition in the conditioning of the merit function at $\lambda=0$, which should be reflected in the number of optimization steps required for convergence. To empirically validate this prediction, we vary both the LLE and sequence length $T$ within a parametric family of recurrent neural networks (RNNs), and measure the number of steps DEER takes to converge. We generate mean-field RNNs following Engelken et al. [29], scaling standard normal weight matrices by a single parameter that controls their variance and therefore the expected LLE. In Figure˜2, we observe a striking correspondence between the conditioning of the optimization problem (represented by $\log\tilde{\mu}$, where $\tilde{\mu}$ is the lower bound for $\mu$ from ˜2) and the number of steps DEER takes to convergence. This relationship holds across the range of LLEs, $\lambda$, and sequence lengths, $T$. There is a rapid threshold phenomenon around $\lambda=0$, which divides predictable from unpredictable dynamics, precisely as expected from ˜2. As we discuss in Appendix J.1, the correspondence between $\log\tilde{\mu}$ and the number of optimization steps needed for convergence can be explained by DEER iterates approaching the basin of quadratic convergence with linear rate.
In Section˜J.3, we provide additional experiments in this setting. We parallelize the sequential rollout with other optimizers like quasi-Newton and gradient descent, and observe that the number of steps these optimizers take to converge also scales with the LLE. We also record wallclock times on an H100, and observe that DEER is faster than sequential by an order of magnitude in predictable settings, but slower by an order of magnitude in unpredictable settings.
@@
To illustrate, we apply DEER to Langevin dynamics in a two-well potential (visualized in Figure˜3 for $D=2$). The dynamics are stable within each well but unstable in the region between them. Despite this local instability, the system’s overall behavior is governed by time spent in the wells, resulting in a negative LLE and sublinear growth in DEER’s convergence steps with sequence length $T$ (Figure 3, right subplot). Additional details and discussion are in Appendix J.4.
@@
Figure 3: DEER converges quickly for Langevin dynamics in a two-well potential. (Left) An illustration of the two-well potential state space in $D=2$. We superimpose a contour plot of the potential on a color scheme showing the spectral norm of the dynamics Jacobian (blue indicates stability, red instability). (Center) A trace plot for the $y$-coordinate. The LLE of the system is 1.45-1.45. (Right) We observe that this system, which has negative LLE, enjoys sublinear scaling in the sequence length $T$ in the number of DEER iterations needed to converge. We plot the median number of DEER steps to convergence over 20 random seeds.
@@
Finally, we demonstrate a practical application of our theory in the efficient parallelization of chaotic observers. Observers are commonly used to reconstruct the full state of a system from partial measurements [30, 31]. On nine chaotic flows from the dysts benchmark dataset [32], Table 1 shows that while DEER converges prohibitively slowly on chaotic systems, it converges rapidly on stable observers of these systems, in accordance with our theory that predictability implies parallelizability. For more details, see Appendix J.5.
Table 1: Comparison of system and observer LLEs and number of DEER steps for $T=30,000$.
@@
6 Conclusion
Recent work has demonstrated that parallel computing hardware like GPUs can be used to rapidly compute state trajectories of nonlinear state space models. The central idea underlying these works is to reconceive the state trajectory as the solution to an optimization problem. Here, we provided a precise characterization of the optimization problem’s inherent difficulty, which determines if parallelization will be faster in practice than sequential evaluation. We show that the conditioning of the optimization problem is governed by the predictability of the underlying nonlinear system. We then translate this insight into worst-case performance guarantees for specific optimizers, including Gauss–Newton (DEER). Our main contribution can be summarized as: Predictable dynamics yield well-conditioned merit functions, enabling rapid convergence. Unpredictable dynamics produce flat or ill-conditioned merit landscapes, resulting in slow convergence or, worse, numerical failure.
@@ 6 Conclusion
The DeepPCR algorithm was introduced by Danieli et al. [2], who investigated its convergence rates empirically but not theoretically. Concurrently, Lim et al. [1] proposed an essentially identical method under the name DEER, proving local quadratic convergence for Gauss–Newton under standard assumptions but leaving the question of global convergence unresolved. Gonzalez et al. [3] proved the global convergence of DEER and other variants, though only with worst-case bounds of $T$ optimization steps. None of these prior works addressed the relationship between system dynamics and conditioning, or established global linear convergence rates.
Global convergence rates for Gauss-Newton are rare, despite the breadth of optimization literature [19, 28, 33]. Theorem 4 establishes global convergence with linear rate for Gauss-Newton by leveraging our specific problem structure.
Fifty years ago, Hyafil and Kung [34] and Kung [35] showed that linear recursions enjoy speedups from parallel processors while nonlinear recursions of rational functions with degree larger than one cannot. These prescient works set the stage for our more general findings, which explicitly link the dynamical properties of the recursion to its parallelizability. Parallel-in-time methods for continuous systems also have a long history [36, 37], with Chartier and Philippe [38] showing that dissipative systems can be parallelized using multiple shooting. Furthermore, Danieli and MacLachlan [39] and De Sterck et al. [40] study the CFL number for determining the usefulness of multigrid systems. Connecting this work with our paper is an interesting direction for future research.
More recently, several works have parallelized diffusion models via fixed-point iteration [2, 11, 41, 42], again with $T$-step worst-case guarantees. Anari et al. [43] proved $\log(T)$ rates for a particular dynamical system (Langevin dynamics) and a particular fixed-point iteration (Picard iteration). Crucially, prior work has not focused on the merit function, which we can define for any discrete-time dynamical system and optimizer.
@@
The centrality of parallel sequence modeling architectures like transformers [44], deep SSMs [45, 21, 46], and linear RNNs [47] in modern machine learning underscores the need for our theoretical work. Merrill et al. [48] explored the question of parallelizability through the lens of circuit complexity, analyzing when deep-learning models can solve structured tasks in constant depth. Their focus complements ours, and suggests an opportunity for synthesis in future work.
@@
Nonlinear state space models are ubiquitous. Our results offer two key contributions.
First, they provide a principled way to determine, a priori, whether optimization-based parallelization of a given model is practical. In many robotic or control systems, particularly ones that are strongly dissipative, this insight can enable orders-of-magnitude speed-ups on GPUs [49, 50, 51, 52, 53, 54, 55, 56].
In concurrent work, Zoltowski et al. [18] developed and leveraged quasi-Newton methods to parallelize Markov Chain Monte Carlo over the sequence length, attaining order-of-magnitude speed-ups. These speed-ups were made possible by the fast convergence of the quasi-Newton methods in the settings considered. Suggestively, there is an abundance of research studying the contractivity of MCMC in different settings [57, 58, 59]. In fact, the empirical results in this paper showing that Langevin dynamics can have negative LLE (cf. Figure˜3) are suggestive that the Metropolis-adjusted Langevin algorithm (MALA), a workhorse of MCMC, may also be predictable in settings of interest. A precise characterization of what makes an MCMC algorithm and target distribution contractive (i.e. predictable, in the language of our paper) would provide useful guidance for when one should aim to parallelize MCMC over the sequence length. Thus, providing precise theoretical justification for parallelizing MCMC over the sequence length is an exciting avenue for future work.
Second, our results have direct implications for system design. When constructing nonlinear dynamical systems in machine learning—such as novel RNNs—parallelization benefits are maximized when the system is made predictable. Given the large body of work on training stable RNNs [60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71], many effective techniques already exist for enforcing stability or predictability during training. A common approach is to parameterize the model’s weights so that the model is always stable. For example, we provide a simple parameterization of a contractive SSM in Appendix˜I.
Notably, the concurrent work of Farsang et al. [71] develops a nonlinear SSM and trains it in parallel using quasi-DEER [3]. Farsang et al. [71] explicitly parameterizes their LrcSSM to be contractive (see their Appendix A.1). While they do not identify it as such, this contractivity is precisely what enables DEER to converge quickly throughout training. Ensuring a negative largest Lyapunov exponent through parameterization guarantees parallelizability for the entire training process, enabling faster and more scalable learning. Our contribution provides a theoretical foundation for why stability is essential in designing efficiently parallelizable nonlinear SSMs.
@@ Acknowledgments
We thank members of the Linderman Lab for helpful feedback, particularly Noah Cowan, Henry Smith, Skyler Wu, and Etaash Katiyar. We thank Federico Danieli for helpful discussions.
X.G. would like to acknowledge support from the Walter Byers Graduate Scholarship from the NCAA. L.K. was a Goldstine Fellow at IBM Research while conducting this research. S.W.L. was supported by fellowships from the Simons Collaboration on the Global Brain, the Alfred P. Sloan Foundation, and the McKnight Foundation.
@@ Appendix B Merit Function is PL
To be precise, we must have $\mu>0$ for $\mathcal{L}$ to satisfy the definition of PL. Therefore, a condition that must apply for $\mathcal{L}$ to be PL is that we must have $\inf_{\mathbf{s}}\sigma_{\min}\left(\mathbf{J}(\mathbf{s})\right)>0$. We note that the proof strategy of ˜2 ensures that $\inf_{\mathbf{s}\in\mathbb{R}^{TD}}\sigma_{\min}\left(\mathbf{J}(\mathbf{s})\right)>0$ if we assume eq.˜21, which holds for dynamical systems that are globally contracting.
By the chain rule, eq.˜21 also holds for functions of the form $f(s)=\phi(Ws)$, where $W\in\mathbb{R}^{D\times D}$ and $\phi$ is a scalar function with bounded derivative that is applied elementwise. In particular, such a function $\phi(Ws)$ satisfies eq.˜21 whether or not it is globally contracting. This function class is extremely common in deep learning (nonlinearities with bounded derivatives include $\tanh$, the logistic function and $\mathrm{ReLU}$).

### v3

@@ Abstract
The rise of parallel computing hardware has made it increasingly important to understand which nonlinear state space models can be efficiently parallelized. Recent advances like DEER [56] and DeepPCR [13] recast sequential evaluation as a parallelizable optimization problem, sometimes yielding dramatic speedups. However, the factors governing the difficulty of these optimization problems remained unclear, limiting broader adoption. In this work, we establish a precise relationship between a system’s dynamics and the conditioning of its corresponding optimization problem, as measured by its Polyak-Łojasiewicz (PL) constant. We show that the predictability of a system, defined as the degree to which small perturbations in state influence future behavior and quantified by the largest Lyapunov exponent (LLE), impacts the number of optimization steps required for evaluation. For predictable systems, the state trajectory can be computed in at worst $\mathcal{O}((\log T)^{2})$ time, where $T$ is the sequence length: a major improvement over the conventional sequential approach. In contrast, chaotic or unpredictable systems exhibit poor conditioning, with the consequence that parallel evaluation converges too slowly to be useful. Importantly, our theoretical analysis shows that predictable systems always yield well-conditioned optimization problems, whereas unpredictable systems lead to severe conditioning degradation. We validate our claims through extensive experiments, providing practical guidance on when nonlinear dynamical systems can be efficiently parallelized. We highlight predictability as a key design principle for parallelizable models.
@@ 1 Introduction
via the Gauss-Newton method. [^1] DEER [56] and DeepPCR [13] were concurrent works that both proposed to use the Gauss-Newton method for optimizing nonlinear sum of squares to parallelize sequential processes. In this paper, we therefore use DEER, DeepPCR, and Gauss-Newton interchangeably. Gonzalez et al. [34] further developed these methods, including quasi-Newton methods and trust-region methods for parallel evaluation of nonlinear dynamical systems. These methods evaluate nonlinear dynamical systems by iteratively linearizing the nonlinear system and evaluating the resulting linear dynamical system (LDS) with a parallel (a.k.a. associative) scan [83, 3]. Each parallel evaluation of an LDS implements one optimization step  [56, 13, 34, 33].
@@ 2 Problem Statement & Background
We define the residual and corresponding merit [^2] While minimizing a “merit function” is admittedly counterintuitive, we follow Nocedal and Wright [69, see eq. 11.35] in this convention. function $\mathcal{L}$ by stacking the elements $s_{t}\in\mathbb{R}^{D}$ of a trajectory into a $TD$-dimensional vector $\mathbf{s}$ and considering the vector of temporal differences,
@@ 3.1 The Merit Function is PL
3.2 Merit Function PL Constant is Controlled by the Largest Lyapunov Exponent of Dynamics
@@ 4 Rates of Convergence for Optimizing the Merit Function
See Appendix H. We make no claim about the originality of lower bounding the size of the basin of quadratic convergence in Gauss-Newton. In fact, our proof of ˜5 closely follows the convergence analysis of Newton’s method in Section 9.5.3 of Boyd and Vandenberghe [5]. Our contribution is we highlight the elegant way the predictability $\lambda$ and nonlinearity $L$ of a dynamical system influence an important feature of its merit function’s landscape. ∎
@@ 5 Experiments
We conduct experiments to support the theory developed above, demonstrating that predictability enables parallelization of nonlinear SSMs. To illustrate this point, we use Gauss-Newton optimization (aka DEER [56, 13]). We provide more experimental details in Appendix K. Our code is at https://github.com/lindermanlab/predictability_enables_parallelization
@@
Theorem 2 predicts a sharp phase transition in the conditioning of the merit function at $\lambda=0$, which should be reflected in the number of optimization steps required for convergence. To empirically validate this prediction, we vary both the LLE and sequence length $T$ within a parametric family of recurrent neural networks (RNNs), and measure the number of steps DEER takes to converge. We generate mean-field RNNs following Engelken et al. [19], scaling standard normal weight matrices by a single parameter that controls their variance and therefore the expected LLE. In Figure˜2, we observe a striking correspondence between the conditioning of the optimization problem (represented by $\log\tilde{\mu}$, where $\tilde{\mu}$ is the lower bound for $\mu$ from ˜2) and the number of steps DEER takes to converge. This relationship holds across the range of LLEs, $\lambda$, and sequence lengths, $T$. There is a rapid threshold phenomenon around $\lambda=0$, which divides predictable from unpredictable dynamics, precisely as expected from ˜2. As we discuss in Appendix K.1, the correspondence between $\log\tilde{\mu}$ and the number of optimization steps needed for convergence can be explained by DEER iterates approaching the basin of quadratic convergence with linear rate.
In Section˜K.3, we provide additional experiments in this setting. We parallelize the sequential rollout with other optimizers like quasi-Newton and gradient descent, and observe that the number of steps these optimizers take to converge also scales with the LLE. We also record wallclock times on an H100, and observe that DEER is faster than sequential by an order of magnitude in predictable settings, but slower by an order of magnitude in unpredictable settings.
@@
To illustrate, we apply DEER to Langevin dynamics in a two-well potential (visualized in Figure˜3 for $D=2$). The dynamics are stable within each well but unstable in the region between them. Despite this local instability, the system’s overall behavior is governed by time spent in the wells, resulting in a negative LLE and sublinear growth in DEER’s convergence steps with sequence length $T$ (Figure 3, right subplot). Additional details and discussion are in Appendix K.4.
@@
Figure 3: DEER converges quickly for Langevin dynamics in a two-well potential. (Left) An illustration of the two-well potential state space in $D=2$. We superimpose a contour plot of the potential on a color scheme showing the spectral norm of the dynamics Jacobian (blue indicates stability, red instability). (Center) A trace plot for the $y$-coordinate. The LLE of the system is 0.0145-0.0145. (Right) We observe that this system, which has negative LLE, enjoys sublinear scaling in the sequence length $T$ in the number of DEER iterations needed to converge. We plot the median number of DEER steps to convergence over 20 random seeds.
@@
Finally, we demonstrate a practical application of our theory in the efficient parallelization of chaotic observers. Observers are commonly used to reconstruct the full state of a system from partial measurements [61, 80]. On nine chaotic flows from the dysts benchmark dataset [29], Table 1 shows that while DEER converges prohibitively slowly on chaotic systems, it converges rapidly on stable observers of these systems, in accordance with our theory that predictability implies parallelizability. For more details, see Appendix K.5.
Table 1: Comparison of system and observer LLEs and number of DEER steps for $T=30,000$ and Euler discretization step size $\Delta t=0.01$.
@@
6 Discussion
Recent work demonstrated that parallel computing hardware like GPUs can be used to rapidly compute state trajectories of nonlinear state space models (nSSMs) by recasting the trajectory as the solution to an optimization problem. In this work, we provide the first precise characterization of the optimization problem’s inherent difficulty, which determines if parallelization will be faster in practice than sequential evaluation. We show that the conditioning of the optimization problem is governed by the predictability of the underlying dynamics. We translate this insight into worst-case performance guarantees for specific optimizers, including Gauss–Newton (DEER). Our main takeaway is: Predictable dynamics yield well-conditioned merit functions, enabling rapid convergence. Unpredictable dynamics produce flat or ill-conditioned merit landscapes, resulting in slow convergence or numerical failure.
@@ 6 Discussion
While Lim et al. [56] and Danieli et al. [13] introduced parallel Newton methods, they did not prove their global convergence. Gonzalez et al. [34] proved global convergence, though only with worst-case bounds of $T$ optimization steps. These prior works did not address the relationship between system dynamics and conditioning, or establish global linear convergence rates.
Global convergence rates for Gauss-Newton are rare, despite the breadth of optimization literature [5, 94]. Theorem 4 establishes global convergence with linear rate for Gauss-Newton by leveraging our specific problem structure, though similar results have existed for local linear convergence [71].
Fifty years ago, Hyafil and Kung [43] and Kung [52] showed that linear recursions enjoy speedups from parallel processors while nonlinear recursions of rational functions with degree larger than one cannot. These prescient works set the stage for our more general findings, which explicitly link the dynamical properties of the recursion to its parallelizability. Parallel-in-time methods for continuous systems also have a long history [68, 26, 70], with Chartier and Philippe [8] showing that dissipative systems can be parallelized using multiple shooting. Furthermore, Danieli and MacLachlan [11] and De Sterck et al. [16] study the CFL number for determining the usefulness of multigrid systems. Connecting this work with our paper is an interesting direction for future research.
More recently, several works have parallelized diffusion models via fixed-point iteration, including worst-case guarantees of $T$ steps [79, 87, 78] as well as polylogarithmic rates in $T$ [1, 9]. Lu et al. [60] develops quasi-Newton methods for sampling from diffusion models and, like us, shows a two-phase model of linear followed by quadratic convergence. Crucially, prior work has not focused on the merit function, which we can define for any discrete-time dynamical system and optimizer.
@@
The centrality of parallel sequence modeling architectures like transformers [89], deep SSMs [37, 82, 36], and linear RNNs [92] in modern machine learning underscores the need for our theoretical work. Merrill et al. [65] explored the question of parallelizability through the lens of circuit complexity, analyzing when deep learning models can solve structured tasks in constant depth. Their focus complements ours, and suggests an opportunity for synthesis in future work [58].
@@
Our work unlocks three key implications for nonlinear state space models.
First, it provides a principled way to determine, a priori, whether optimization-based parallelization of a given model is practical. In many robotic or control systems, particularly ones that are strongly dissipative, this insight can enable orders-of-magnitude speed-ups on GPUs [49, 2, 44, 22, 81, 85, 88, 75, 15].
For example, the concurrent work of Zoltowski et al. [95] developed and leveraged quasi-Newton methods to parallelize Markov Chain Monte Carlo over the sequence length, attaining order-of-magnitude speed-ups. These speed-ups occurred because the quasi-Newton methods converged quickly in the settings considered. Suggestively, MCMC chains are contractive in many settings [4, 63, 18]. A precise characterization of what makes an MCMC algorithm and target distribution predictable would provide useful guidance for when one should aim to parallelize MCMC over the sequence length. Providing precise theoretical justification for parallelizing MCMC over the sequence length is an exciting avenue for future work.
Second, our results impact architecture design. When constructing nonlinear dynamical systems in machine learning—such as novel RNNs—parallelization benefits are maximized when the system is made predictable. Given the large body of work on training stable RNNs [41, 40, 86, 66, 21, 50, 32, 51, 20, 72, 96, 23], many effective techniques already exist for enforcing stability or predictability during training. A common approach is to parameterize the model’s weights so that the model is always stable (see Appendix˜I).
Notably, the concurrent work of Farsang et al. [23] and Danieli et al. [12] develop nonlinear SSMs and train them with DEER, with Danieli et al. [12] scaling to very strong performance as a 7B parameter language model. Both highlight the fast convergence of DEER, which is a result of the contractivity of their architectures: Farsang et al. [23] parameterizes their LrcSSM to be contractive, while Danieli et al. [12] clip the norms of their weight matrices. Ensuring a negative largest Lyapunov exponent through parameterization guarantees parallelizability for the entire training process, enabling faster and more scalable learning. Our contribution provides a theoretical foundation for why stability is essential in designing efficiently parallelizable nonlinear SSMs.
Finally, our results have implications for the interpretation of stable nSSMs. Because each Gauss-Newton step in DEER is a linear dynamical system (LDS), and because we prove in ˜4 that DEER converges in $\mathcal{O}(\log T)$ steps for a stable nSSM, we can interpret a stable nSSM as being equivalent to a “stack” of $\mathcal{O}(\log T)$ LDSs coupled by nonlinearities (cf. Appendix˜J).
@@ Acknowledgments
We thank members of the Linderman Lab for helpful feedback, particularly Noah Cowan, Henry Smith, Skyler Wu, and Etaash Katiyar. We also thank Federico Danieli, Will Merrill, Julien Siems, and Riccardo Grazzi for helpful discussions. We thank the anonymous NeurIPS reviewers whose feedback improved this paper.
X.G. acknowledges support from the Walter Byers Graduate Scholarship from the NCAA. L.K. was a Goldstine Fellow at IBM Research while conducting this research. S.W.L. was supported by fellowships from the Simons Collaboration on the Global Brain, the Alfred P. Sloan Foundation, and the McKnight Foundation.
@@ Appendix B Merit Function is PL
To be precise, we must have $\mu>0$ for $\mathcal{L}$ to satisfy the definition of PL. Therefore, a condition that must apply for $\mathcal{L}$ to be PL is that we must have $\inf_{\mathbf{s}}\sigma_{\min}\left(\mathbf{J}(\mathbf{s})\right)>0$. We note that the proof strategy of ˜2 ensures that $\inf_{\mathbf{s}\in\mathbb{R}^{TD}}\sigma_{\min}\left(\mathbf{J}(\mathbf{s})\right)>0$ if we assume eq.˜22, which holds for dynamical systems that are globally contracting.
By the chain rule, eq.˜22 also holds for functions of the form $f(s)=\phi(Ws)$, where $W\in\mathbb{R}^{D\times D}$ and $\phi$ is a scalar function with bounded derivative that is applied elementwise. In particular, such a function $\phi(Ws)$ satisfies eq.˜22 whether or not it is globally contracting. This function class is extremely common in deep learning (nonlinearities with bounded derivatives include $\tanh$, the logistic function and $\mathrm{ReLU}$).